DIRC Operations Report
Period from 2008/02/19 to 2008/03/19
Oncall: Jennifer Prendki, Jose Benitez
Fast Monitoring: Jennifer Prendki, Jose Benitez
PEP peak lumi: ~12.0 x 1033
Run 7 total delivery: ~45.90 fb-1
~0.62 fb-1 on the Y(4S),
~30.66 fb-1 on the Y(3S),
~12.12 fb-1 on the Y(2S).
PEP, BaBar activities/issues since last DIRC meeting
Took some Y(3S) offpeak data.
Switched to running on the Y(2S) on March 1st.
Backgrounds are high since start of Y(2S) data.
Took some Y(2S) offpeak data on Tuesday.
DIRC activities/issues since last DIRC meeting
Shoulder in timing distributions was back on Monday Swing and Tuesday Day shifts. Went worse when switching to Y(2S) offpeak. Fast monitoring plots showed trigger lines #1 & #8 were responsible for that.
--> Things are now back to normal.
Current detector status :
We have a new broken chip in Sector #8.
One chip in Sector #0 seems to be dying those days. We keep an eye on this.
See DIRC logbook for details
Lumi and Background History
Some more plots for background history :
The following plot shows the correlation between DIRC scaler and the Neutron counter ...
... And those ones the correlations bewteen the DIRC scalers and LST backgrounds in inner layer.
The LSTs have been struggling with backgrounds. Since PEP had nothing to tune against (until the new stripcharts were added), decision was made to use the scaler DIRC alarms as a reference for them. Therefore the alarm threshold was not increase even though those persistent alarms have been bothering pilot shifters for a while in IR2.
| Equipment || # used || Spares |
| DFB || 168
12/17 good + 4 possibly good, 1 bad
| DCC || 12
| HV Boards || 48
9/11 good , 2 broken recovered by Jose from Angels shop because they couldn't fix them
- I (Jennifer) am carrying the pager, Jose is in backup.
- Thanks once again to Joe for the great job !